OpenStudy (anonymous):

how do I simplify sqrt(1-cos^2(2x))? I think I should get sin(2x)

OpenStudy (dumbcow):

yes that is correct

OpenStudy (anonymous):

oh wait do I just sub in sin^2 for 1-cos^2?

OpenStudy (dumbcow):

yep...pythagorean identity

OpenStudy (anonymous):

Say we have \(y=\sqrt{1-cos^2(2x)}\) Squaring both sides gives: \(y^2=1-cos^2(2x)\) So \(y^2 + cos^2(2x)=1\) We know that for any angle, \(sin^2+cos^2=1\). So y is clearly \(sin(2x)\). Does that explanation make sense?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

No worries :)